OpenStudy (anonymous):

which of the following functions is not odd? a. f(X)= sinx b. f(X) = sin2x c. f(X)= x^3 +1 d. f(X)= x/(x^2+1) e. f(X)= 3rdroot2x please show how to find solution!!

OpenStudy (anonymous):

c. is not odd because f(-x) = (-x)^3+1= -x^3+1 does NOT equal -f(x) = -x^3-1.